Suppose a layer of oil (refractive index is 1.45) floats on water (refractive index is 1.33).If a ray of light shines onto the oil with an incidence angle of 40°, what is the angle formes between the ray and the surface of the water?





a) Given;



b)Formula:



c)Solution



d)Final answer with unit:

1
Expert's answer
2022-05-04T17:38:46-0400

"\alpha=90°-40°=50°,"

"\frac{\sin\alpha}{\sin\beta}=n_1,"

"\beta=31.9°,"

"\beta'=90°-\beta=58.1°."